RRRG new website top logo resized

 

Restoration diary

Reports of restoration work on 50029 and 50030.

Announcement of working party dates and details of restoration work carried out at working parties.

Working party report 28th June 2015

Working party report 28th June 2015
This work party was attended by myself plus the "Three Amigos": Dave Rolfe, Ian Kemp and Peter Carter.  Dave and I fitted a couple more water and oil pipes to the ex-50008 power unit, I then swapped a water pipe from the power unit inside 50030 to replace a damaged one on the new power unit.
 
The main news from this Sunday centres around the heat exchanger which was removed from storage, due to the size and weight this was no mean feat but between the four of us and the use of our winch we managed to move it out for inspection. When we purchased the ex-50008 power unit we were told the heat exchanger had been swapped for the one currently in 50008 as it had a leak. When Ian inspected it, it was clear the end plate had corroded through in a couple of places (this plate is around 5mm thick so it's surprising it has rusted so badly). Thankfully the rest of the exchanger looks to be ok but we will need to pressure test it at some point. The rest of the afternoon was spent removing the end plate from the heat exchanger inside 50030; although this is quite rusty it appears to be in much better condition and should be easy(ish) to swap over.

 

 b2ap3_thumbnail_oil-pipe-and-gasket-med.jpg b2ap3_thumbnail_heat-exchanger-med.jpg 
Continue reading
3719 Hits
0 Comments

Working party report 31st May 2015

Working party report 31st May 2015

Ian Kemp, Peter Carter and Dave Rolfe carried on fitting oil, water and fuel pipework to the ex-50008 power unit, but on the side facing Renown, so no photo opportunities were possible today.

In preparation for the open day on June 21st, I washed down the outside of the electrical cubicle frame in 50030 and undercoated it, as it had become dirty from repeated handling since it was originally painted about five or six years ago. I also cleaned and undercoated the two cubicle doors from the fuse and MCB compartment. I brought three more cubicle doors home to clean.

b2ap3_thumbnail_010.JPG b2ap3_thumbnail_008.JPG 

Just a reminder that we are having our Open Day on the afternoon of Sunday June 21st where RRRG members will be on hand to give guided tours of the project and discuss progress. We hope to see you there!

Continue reading
3369 Hits
0 Comments

Working party report 24th May 2015

Working party report 24th May 2015

Ian Kemp, Peter Carter and Wes Needle continued fitting coolant and oil pipework to the power unit, while Dave Rolfe carried on painting the other side of the power unit.

a1sx2_Display_008.JPG

I removed the battery charge contactor so I could clean and undercoat the compartment where the CU2 and CU7 go, then did the same for the compartment underneath where the CU1 sits, in readiness for having them temporarily fitted for the open day. Topcoat will be applied next workparty. I also cleaned the dirtiest two of the cubicle doors ready for painting.

a1sx2_Display_005-2.JPG a1sx2_Display_006-2.JPG

Continue reading
3339 Hits
0 Comments

Working party report 17th May 2015

Mark Burrows and myself were on site today.  A slight mix-up over keys to the container resulted in a delay getting started and curtailed the working day somewhat.  Mark scraped off the worst of the crap from both the 037 cab and no. 2 cab of 30, then painted grey undercoat on the cab front of 037 to 'smarten it up a bit'.  The paintwork on all of our assets is looking faded and dirty after several years of no attention.  I pulled out the cables from under the floor in front of the aux gen that go to the load limiting potentiometer to bell them out. I identified and marked all but one, which will require further investigation. I did the same for several smaller severed cables from the cubicle to the main gen. This took all day as I was doing it on my own, so I had to make repeated trips into the Thin Man's passage to move the belling lead.  Although the work I've described here doesn't sound like much, it did take up a whole day but was very satisfying as it pinned down more of the cabling inside 50030.

 

b2ap3_thumbnail_004-2.JPG b2ap3_thumbnail_005.JPG

Continue reading
3199 Hits
0 Comments

Working party report 19th April 2015

Working party report 19th April 2015

During the work party of 19th April Andy Rowlands connected up the reverser air piston coils and tidied up the cables in the electrical cubicle in 50030, then started investigating the severed cables under the main cubicle walkway. They appear to be from the DSD relay box but further investigation will be required to trace them exactly.

b2ap3_thumbnail_004.JPG b2ap3_thumbnail_006.JPG

Peter Carter and Ian Kemp together with Dave Rolfe continued fitting oil and water pipes to the ex-50008 power unit (see previous report, Power unit progress) in continued preparation for its use inside the overhauled 50030.

b2ap3_thumbnail_009.JPG b2ap3_thumbnail_010.JPG

 

Continue reading
3603 Hits
0 Comments